### v1

@@ Abstract
In clinical practice, medical image segmentation provides useful information on the contours and dimensions of target organs or tissues, facilitating improved diagnosis, analysis, and treatment. In the past few years, convolutional neural networks (CNNs) and Transformers have dominated this area, but they still suffer from either limited receptive fields or costly long-range modeling. Mamba, a State Space Sequence Model (SSM), recently emerged as a promising paradigm for long-range dependency modeling with linear complexity. In this paper, we introduce a Large Window-based Mamba U-shape Network, or LMa-UNet, for 2D and 3D medical image segmentation. A distinguishing feature of our LMa-UNet is its utilization of large windows, excelling in locally spatial modeling compared to small kernel-based CNNs and small window-based Transformers, while maintaining superior efficiency in global modeling compared to self-attention with quadratic complexity. Additionally, we design a novel hierarchical and bidirectional Mamba block to further enhance the global and neighborhood spatial modeling capability of Mamba. Comprehensive experiments demonstrate the effectiveness and efficiency of our method and the feasibility of using large window size to achieve large receptive fields. Codes are available at https://github.com/wjh892521292/LMa-UNet.
@@ 1 Introduction
In this paper, we propose a Large-window-based Mamba U-shape Network (LMa-UNet) for 2D and 3D medical image segmentation. LMa-UNet utilizes the powerful sequence modeling capabilities and linear complexity of Mamba to implement large receptive fields by assigning large windows to SSM modules. Further, we design a novel hierarchical and bidirectional large window-based Mamba block (LM block) to enhance the representation modeling capability of SSMs. The bidirectional Mamba design is used for location-aware sequence modeling, reducing the weight impact of the input order. The hierarchical Mamba design is composed of two types of SSM operations: (i) pixel-level SSM (PiM) and (ii) patch-level SSM (PaM). PiM captures the neighborhood and local-scope pixel information by large window SSM and avoids the forgetting problem that occurs in SSMs due to discontinuity of adjacent patches after tokenization. PaM deals with long-range dependency modeling and global patch interaction.
Our main contributions are as follows. (1) We propose a Large window-based Mamba UNet (LMa-UNet) for 2D/3D medical image segmentation. (2) We assign windows of large receptive fields to SSM layers to enable the model to possess the capability of large spatial modeling. (3) We design a bidirectional Mamba for location-aware sequence modeling. (4) We propose a novel hierarchical Mamba module composed of pixel-level SSM (PiM) and patch-level SSM (PaM), enhancing local-neighborhood pixel-level feature modeling and long-range global patch-level modeling.
@@ 2 Preliminaries: SSM Models
After the discretization of $\mathbf{A},\mathbf{B}$, the discretized version of Eq. (1) (i.e., SSM models) can be written as:
@@ 3 Method
In this section, we first introduce the overall architecture of LMa-UNet. Subsequently, we elaborate the core component, the LM block.
Figure 1: An overview of our proposed LMa-UNet.
3.1 LMa-UNet
An overview of LMa-UNet is given in Fig. 1. Specifically, except for the common UNet composition of a depth-wise convolution, an encoder with downsampling layers, a decoder with upsampling layers, and skip-connection, LMa-UNet improves the structure of UNet with proposed large window-based Mamba (LM) blocks inserted into the encoder. Given a 3D input image with a resolution of $C\times D\times H\times W$, the depth-wise convolution first encodes the input into a feature map $\mathbf{F}_{0}\in\mathcal{R}^{48\times\frac{D}{2}\times\frac{H}{2}\times\frac{W}{2}}$. Then the feature map $\mathbf{F}_{0}$ is fed into each LM block and the corresponding down-sampling layers, and multi-scale feature maps are obtained. An LM block contains two Mamba modules: pixel-level SSM (PiM) and patch-level SSM (PaM). For the $l^{th}$ layer, the process can be formulated as:
@@ 3.2 LM block
LM block is our core component used for further spatial modeling of feature maps of different scales at each stage. Different from the previous methods that use CNN for local pixel-level modeling and Transformer for long-range patch-level dependency modeling, an LM block can accomplish pixel-level and patch-level modeling simultaneously, benefiting from the linear complexity of Mamba. More critically, a lower complexity allows setting larger windows to obtain bigger receptive fields, which will improve the efficiency of local modeling, as shown in Fig. 2(a). Specifically, LM block is a hierarchical design that consists of pixel-level SSM (PiM) and patch-level SSM (PaM); the former is used for local-neighborhood pixel modeling and the latter is used for global long-range dependency modeling. Besides, each Mamba layer in an LM block is bidirectional, which is proposed for location-aware sequence modeling.
@@
Pixel-level SSM (PiM). Since Mamba is a continuous model, the discrete nature of input pixels can weaken the correlation modeling of locally adjacent pixels. Hence, we propose a pixel-level SSM to split the feature map into multiple large sub-windows and perform SSM operations on the sub-windows. We first equally divide a whole feature map into non-overlapping sub-windows for 2D or sub-cubes for 3D. Take 2D for example. Given an input of $H\times W$ resolution, we divide the feature map into sub-windows of size $m\times n$ each ($m$ and $n$ can be up to 40). Without loss of generality, we assume that $H/m$ and $W/n$ are both integers. Then we have $\frac{HW}{mn}$ sub-windows, as shown in Pixel-level SSM of Fig. 1. Under this scheme, when these sub-windows are sent into a Mamba layer, the local adjacent pixels will be input continuously into SSM; thus, the relation between local-neighborhood pixels can be better modeled. Moreover, under the large window partition strategy, the receptive fields are enlarged and the model can obtain more details of the local pixels. However, the image is divided into non-overlapping sub-windows. Hence, we need a mechanism for communication between different sub-windows, for long-range dependency modeling.
@@
Figure 2: (a) Respective field comparison among CNN, Transformer, and our proposed LMa-UNet. CNNs often use small kernels (like $3\times 3$), and Transformers often use $7\times 7$ sized windows. (b) Scanning order comparison of vanilla Mamba vs. our proposed bidirectional Mamba.
@@
Patch-level SSM (PaM). We introduce a patch-level SSM layer to impart information among different sub-windows. As shown in Patch-level SSM of Fig. 1, a feature map $\mathbf{F^{\prime}}_{l}$ of resolution $H\times W$ first passes through a pooling layer of size $m\times n$ to allow important information for each of the $\frac{HW}{mn}$ sub-windows to be summarized into a single representative. Thus, we obtain aggregate maps $\mathbf{Z}_{l}$ with $\frac{HW}{mn}$ representatives, which are then used to communicate among the sub-windows through Mamba for global-range dependency modeling. After the interaction in Mamba, we unpooling the aggregate maps back to the same size as the initial feature map $\mathbf{F^{\prime}}_{l}$, and apply a residual connection. The process of $\mathbf{F}^{\prime\prime}_{l}=\text{PaM}(\mathbf{F}^{\prime}_{l})$ in Eq. (5) can be carried out as:
@@ 4.1 Datasets
We conduct experimental comparisons with state-of-the-art methods on two datasets for 2D and 3D segmentation tasks to validate the effectiveness and scalability of LMa-UNet.
@@ 4.2 Implementation Setup
Our LMa-UNet is implemented on PyTorch 1.9.0 based on the nnU-Net framework. All the experiments are conducted on an NVIDIA GeForce RTX 3090 GPU. The batch size in training is 2 for the 3D dataset (Abdomen CT) and 24 for the 2D dataset (Abdomen MR). The Adam [12] optimizer with momentum = 0.99 is used. The initial learning rate is 0.01 with a weight decay of 3e-5. The maximum training epoch number is 1000. For the Abdomen CT dataset, the stage is 6 but the dimensions are not consistent; thus we set the rectangle window size with three dimensions to [20, 28, 24], [20, 28, 24], [10, 14, 12], [10, 14, 12], [5, 7, 6], and [5, 7, 6] for each stage. For the Abdomen MR dataset, the stage is 7 and the window size is 40, 20, 20, 10, 10, 5, and 5 for each stage.
@@ 4.3 Overall Performances
The baseline models include three types of representative medical image segmentation networks: CNN-based networks (nnU-Net [10] and SegResNet [17]), Transformer-based networks (UNETR [8], SwinUNETR [7], and nnFormer [24]), and the latest Mamba-based network (U-Mamba [15]). For fair comparison, we also implement all the models in the nnU-Net framework, and use the default image pre-processing. Table 1 presents the results. Compared to both CNN-based and Transformer-based segmentation methods, our proposed LMa-UNet achieves improved performances in both DSC and NSD, which indicates that the global modeling capabilities of Mamba are critical to medical image segmentation. Note that, compared to U-Mamba which simply applies Mamba as a global modeling adapter, LMa-UNet exhibits improvements over U-Mamba, validating the effectiveness of our bidirectional and hierarchical Mamba designs. These results also demonstrate the potential of Mamba in global and local feature modeling with large window-based receptive fields.
Table 2: Performances of LMa-UNet in three different window size settings. The window size sequence indicates the window size in each stage.
Window size [10, 5, 5, 5, 5, 5, 5] [20, 10, 10, 10, 5, 5, 5] [40, 20, 20, 10, 10, 5, 5]
@@
Table 3: Performances of LMa-UNet with different sub-modules. PiM = Pixel-level SSM. PaM = Patch-level SSM. BiM = Bidirectional Mamba.
@@
4.4 Is the Window Size of LMa-UNet Important?
Due to the linear complexity of Mamba, LMa-UNet can achieve large window modeling that CNNs and Transformers cannot. To validate Mamba’s large spatial modeling potential, we explore LMa-UNet’s performance in different window size settings. Table 2 shows the results on the Abdomen MR dataset, whose stage number is 7. Comparing the performances of the three window-size settings, one can see that LMa-UNet with larger window sizes achieves better performances. This indicates that large receptive fields are critical for medical image segmentation which can be achieved with Mamba due to its linear complexity. This will provide more possibilities for future large kernel or window modeling.
@@ 4.5 Ablation Study
To validate the effect of each key component in our LMa-UNet, we conduct ablation experiments on the Abdomen MR dataset. Table 3 shows the results. One can see that both PiM and PaM provide improvements for LMa-UNet over the baseline model, validating the effectiveness of PiM and PaM in local pixel-level modeling and global modeling, respectively. Notably, the model with PiM gains more improvements than that with PaM, suggesting that enlarging the receptive field of local feature modeling is a key to improving model performance. After introducing BiM, the performance of LMa-UNet further improves, which shows the importance of bidirectional Mamba for location-aware sequence modeling. Finally, LMa-UNet with all the components achieves the best performance, which further demonstrates the effectiveness of our method and its components.
@@ 5 Conclusions
In this paper, we introduced a new Mamba-based UNet model for medical image segmentation, achieving large window spatial modeling. Further, we designed a bidirectional and hierarchical SSM to enhance the capacities of Mamba in local and global feature modeling. Comprehensive experiments on multi-organ segmentation datasets demonstrated the effectiveness of our proposed method. In the future, we will explore the extensibility and applications of SSMs in other medical imaging tasks, such as detection, registration, reconstruction, etc.

### v2

@@ Abstract
In clinical practice, medical image segmentation provides useful information on the contours and dimensions of target organs or tissues, facilitating improved diagnosis, analysis, and treatment. In the past few years, convolutional neural networks (CNNs) and Transformers have dominated this area, but they still suffer from either limited receptive fields or costly long-range modeling. Mamba, a State Space Sequence Model (SSM), recently emerged as a promising paradigm for long-range dependency modeling with linear complexity. In this paper, we introduce a Large Kernel vision Mamba U-shape Network, or LKM-UNet, for medical image segmentation. A distinguishing feature of our LKM-UNet is its utilization of large Mamba kernels, excelling in locally spatial modeling compared to small kernel-based CNNs and Transformers, while maintaining superior efficiency in global modeling compared to self-attention with quadratic complexity. Additionally, we design a novel hierarchical and bidirectional Mamba block to further enhance Mamba’s global and neighborhood spatial modeling capability for vision inputs. Comprehensive experiments demonstrate the feasibility and the effectiveness of using large-size Mamba kernels to achieve large receptive fields. Codes are available at https://github.com/wjh892521292/LKM-UNet.
@@ 1 Introduction
In this paper, we propose a Large Kernel Mamba U-shape Network (LKM -UNet) for 2D and 3D medical image segmentation. LKM-UNet utilizes the powerful sequence modeling capabilities and linear complexity of Mamba to implement large receptive fields by assigning large kernels (windows) to SSM modules. Further, we design a novel hierarchical and bidirectional large kernel Mamba block (LM block) to enhance the representation modeling capability of SSMs. The bidirectional Mamba design is used for location-aware sequence modeling, reducing the weight impact of the input order. The hierarchical Mamba design is composed of two types of SSM operations: (i) pixel-level SSM (PiM) and (ii) patch-level SSM (PaM). PiM captures the neighborhood and local-scope pixel information by large kernel SSM and avoids the forgetting problem that occurs in SSMs due to discontinuity of adjacent patches after tokenization. PaM deals with long-range dependency modeling and global patch interaction.
Our main contributions are as follows. (1) We propose a Large Kernel Mamba UNet (LKM-UNet) for 2D/3D medical image segmentation. (2) We assign kernels of large receptive fields to SSM layers to enable the model to possess the capability of large spatial modeling. (3) We design a bidirectional Mamba for location-aware sequence modeling. (4) We propose a novel hierarchical Mamba module composed of pixel-level SSM (PiM) and patch-level SSM (PaM), enhancing local-neighborhood pixel-level and long-range global patch-level modeling.
@@ 2 Preliminaries: SSM Models
After the discretization, the discretized version of Eq. (1) can be defined as :
@@ 3 Method
In this section, we first introduce the overall architecture of LKM-UNet. Subsequently, we elaborate the core component, the LM block.
Figure 1: An overview of our proposed LKM-UNet.
3.1 LKM-UNet
An overview of LKM-UNet is given in Fig. 1. Specifically, except for the common UNet composition of a depth-wise convolution, an encoder with downsampling layers, a decoder with upsampling layers, and skip-connection, LKM-UNet improves the structure of UNet with proposed large kernel Mamba (LM) blocks inserted into the encoder. Given a 3D input image with a resolution of $C\times D\times H\times W$, the depth-wise convolution first encodes the input into a feature map $\mathbf{F}_{0}\in\mathcal{R}^{48\times\frac{D}{2}\times\frac{H}{2}\times\frac{W}{2}}$. Then the feature map $\mathbf{F}_{0}$ is fed into each LM block and the corresponding down-sampling layers, and multi-scale feature maps are obtained. An LM block contains two Mamba modules: pixel-level SSM (PiM) and patch-level SSM (PaM). For the $l^{th}$ layer, the process can be formulated as:
@@ 3.2 LM block
LM block is our core component used for further spatial modeling of feature maps of different scales at each stage. Different from the previous methods that use CNN for local pixel-level modeling and Transformer for long-range patch-level dependency modeling, an LM block can accomplish pixel-level and patch-level modeling simultaneously, benefiting from the linear complexity of Mamba. More critically, a lower complexity allows setting larger kernels (windows) to obtain bigger receptive fields, which will improve the efficiency of local modeling, as shown in Fig. 2(a). Specifically, LM block is a hierarchical design that consists of pixel-level SSM (PiM) and patch-level SSM (PaM); the former is used for local-neighborhood pixel modeling and the latter is used for global long-range dependency modeling. Besides, each Mamba layer in an LM block is bidirectional, which is proposed for location-aware sequence modeling.
@@
Pixel-level SSM (PiM). Since Mamba is a continuous model, the discrete nature of input pixels can weaken the correlation modeling of locally adjacent pixels. Hence, we propose a pixel-level SSM to split the feature map into multiple large sub-kernels (sub-windows) and perform SSM operations on the sub-kernels. We first equally divide a whole feature map into non-overlapping sub-kernels for 2D or sub-cubes for 3D. Take 2D for example. Given an input of $H\times W$ resolution, we divide the feature map into sub-kernels of size $m\times n$ each ($m$ and $n$ can be up to 40). Without loss of generality, we assume that $H/m$ and $W/n$ are both integers. Then we have $\frac{HW}{mn}$ sub-kernels, as shown in Pixel-level SSM of Fig. 1. Under this scheme, when these sub-kernels are sent into a Mamba layer, the local adjacent pixels will be input continuously into SSM; thus, the relation between local-neighborhood pixels can be better modeled. Moreover, under the large kernel partition strategy, the receptive fields are enlarged and the model can obtain more details of the local pixels. However, the image is divided into non-overlapping sub-kernels. Hence, we need a mechanism for communication between different sub-kernels, for long-range dependency modeling.
@@
Figure 2: (a) Respective field comparison among CNN, Transformer, and our proposed LKM-UNet. CNNs often use small kernels (like $3\times 3$), and Transformers often use $7\times 7$ sized kernels (windows). Our LKM-UNet can scale up kernel size to $40\times 40$. (b) Scanning order comparison of vanilla Mamba vs. our proposed bidirectional Mamba.
@@
Patch-level SSM (PaM). We introduce a patch-level SSM layer to impart information among different sub-kernels. As shown in Patch-level SSM of Fig. 1, a feature map $\mathbf{F^{\prime}}_{l}$ of resolution $H\times W$ first passes through a pooling layer of size $m\times n$ to allow important information for each of the $\frac{HW}{mn}$ sub-kernels to be summarized into a single representative. Thus, we obtain aggregate maps $\mathbf{Z}_{l}$ with $\frac{HW}{mn}$ representatives, which are then used to communicate among the sub-kernels through Mamba for global-range dependency modeling. After the interaction in Mamba, we unpooling the aggregate maps back to the same size as the initial feature map $\mathbf{F^{\prime}}_{l}$, and apply a residual connection. The process of $\mathbf{F}^{\prime\prime}_{l}=\text{PaM}(\mathbf{F}^{\prime}_{l})$ in Eq. (5) can be carried out as:
@@ 4.1 Datasets
We conduct experimental comparisons with state-of-the-art methods on two datasets for 2D and 3D segmentation tasks to validate the effectiveness and scalability of LKM-UNet.
@@ 4.2 Implementation Setup
Our LKM-UNet is implemented on PyTorch 1.9.0 based on the nnU-Net framework. All the experiments are conducted on an NVIDIA GeForce RTX 3090 GPU. The batch size in training is 2 for the 3D dataset (Abdomen CT) and 24 for the 2D dataset (Abdomen MR). The Adam [12] optimizer with momentum = 0.99 is used. The initial learning rate is 0.01 with a weight decay of 3e-5. The maximum training epoch number is 1000. For the Abdomen CT dataset, the stage is 6 but the dimensions are not consistent; thus we set the rectangle kernel size with three dimensions to [20, 28, 24], [20, 28, 24], [10, 14, 12], [10, 14, 12], [5, 7, 6], and [5, 7, 6] for each stage. For the Abdomen MR dataset, the stage is 7 and the kernel size is 40, 20, 20, 10, 10, 5, and 5 for each stage.
@@ 4.3 Overall Performances
The baseline models include three types of representative networks: CNN-based networks (nnU-Net [10] and SegResNet [17]), Transformer-based networks (UNETR [8], SwinUNETR [7], nnFormer [24]), and the latest Mamba-based network (U-Mamba [15]). For fair comparison, we also implement all the models in the nnU-Net framework, and use the default image pre-processing. Table 1 presents the results. Compared to both CNN-based and Transformer-based segmentation methods, our proposed LKM-UNet achieves improved performances in both DSC and NSD, which indicates that the global modeling capabilities of Mamba are critical to medical image segmentation. Note that, compared to U-Mamba which simply applies Mamba as a global modeling adapter, LKM-UNet exhibits improvements over U-Mamba, validating the effectiveness of our bidirectional and hierarchical Mamba designs. These results also demonstrate the potential of Mamba in global and local feature modeling with larger receptive fields.
Table 2: Performances of LKM-UNet in three different kernel size settings. The kernel size sequence indicates the kernel size in each stage (the total number of stages is 7).
Kernel size [10, 5, 5, 5, 5, 5, 5] [20, 10, 10, 10, 5, 5, 5] [40, 20, 20, 10, 10, 5, 5]
@@
Table 3: Performances of LKM-UNet with different sub-modules. PiM = Pixel-level SSM. PaM = Patch-level SSM. BiM = Bidirectional Mamba.
@@
4.4 Is the Kernel Size of LKM-UNet Important?
To validate Mamba’s large spatial modeling potential, we explore LKM-UNet’s performance in different kernel size settings. Table 2 shows the results on the Abdomen MR dataset. Comparing the performances of the three kernel-size settings, one can see that LKM-UNet with larger kernel sizes achieves better performances. This indicates that large receptive fields are critical for medical image segmentation which can be achieved with Mamba due to its linear complexity.
@@ 4.5 Ablation Study
We conduct ablation experiments on the Abdomen MR dataset to validate the effect of each key component in our LKM-UNet, shown in Table 3. Both PiM and PaM provide improvements for LKM-UNet over the baseline model, validating the superiority of PiM and PaM in local pixel-level modeling and global modeling, respectively. Notably, the model with PiM gains more improvements, suggesting that enlarging the receptive field of local modeling is a key to improving model performance. After introducing BiM, the performance of LKM-UNet further improves, which shows the importance of bidirectional Mamba for location-aware modeling. Finally, LKM-UNet with all the components achieves the best performance, further demonstrating our method’s effectiveness and its components.
4.6 Effective Receptive Field Visualization
To show more details of receptive field, we exhibit the effective receptive field (ERF) [14] of other methods and LKM-UNet in Fig. 3. CNN-based methods focus more on local feature extraction, while Transformer-based methods have a wider range of ERF. Although U-Mamba utilize Mamba to obtain globally ERF, it weakens some local attentions. By contrast, our proposed LKM-UNet with large kernel Mamba achieves larger ERF both in global and local aspects.
Figure 3: Effective respective field visualization among CNN, Transformer, U-Mamba and our proposed LKM-UNet.
@@ 5 Conclusions
In this paper, we introduced a new Mamba-based UNet model for medical image segmentation, achieving large kernel spatial modeling. Further, we designed a bidirectional and hierarchical SSM to enhance the capacities of Mamba in local and global feature modeling. Comprehensive experiments on multi-organ segmentation datasets demonstrated the effectiveness of our proposed method.
{credits}
5.0.1 Acknowledgements
This research was partially supported by National Natural Science Foundation of China under grants No. 62176231, No. 62106218, No. 82202984, No. 92259202 and No. 62132017, Zhejiang Key R&D Program of China under grant No. 2023C03053.111The authors have no competing interests to declare that are relevant to the content of this article.